OpenStudy (harrycraig):

If one root of the equation 8x^2-6x-a-3=0 is the square of the other , then the value of a ?

OpenStudy (harrycraig):

@jim_thompson5910

OpenStudy (harrycraig):

\[8x^2-6x-a-3\]

jimthompson5910 (jim_thompson5910):

Hint: one root is r, so the other root is r^2 (x - r)*(x - r^2) = x^2 - r^2*x - r*x + r^3

OpenStudy (sshayer):

let the roots be t and t^2 \[\sum~of~roots=t+t^2=\frac{ -b }{ a }=-\frac{ -6 }{ 8 }=\frac{ 3 }{ 4 }\] product of roots=t*t^2=c/a=\[\frac{ -a-3 }{ 8 }\] \[t^3=\frac{ -a-3 }{ 8 }\] \[t+t^2=\frac{ 3 }{ 4 }\] cubing both sides \[\left( t+t^2 \right)^3=\left( \frac{ 3 }{ 4 } \right)^3\] \[t^3+t^6+3t*t^2\left( t+t^2 \right)=\frac{ 27 }{ 64 }\] \[\frac{ -a-3 }{ 8 }+\left( \frac{ -a-3 }{ 8 } \right)^2+3*\frac{ -a-3 }{ 8 }\left( \frac{ 3 }{ 4 } \right)=\frac{ 27 }{ 64 }\] multiply by 64 \[-8a-24+a^2+9+6a-18a-54=27,a^2-20a-72=0,a=\frac{ 20\pm \sqrt{400+288} }{ 2 }\] a=?

OpenStudy (harrycraig):

If one root is square of other root . question is this

OpenStudy (sshayer):

yes in \[ax^2+bx+c=0\] sum of roots=-b/a product of roots=c/a

OpenStudy (harrycraig):

i don't know but your answer isn't correct

OpenStudy (harrycraig):

@jim_thompson5910 help

OpenStudy (harrycraig):

@sshayer you did mistake on first step

jimthompson5910 (jim_thompson5910):

(x - r)*(x - r^2) = x^2 - r^2*x - r*x + r^3 8*(x - r)*(x - r^2) = 8*(x^2 - r^2*x - r*x + r^3) 8*(x - r)*(x - r^2) = 8*x^2 + 8*( -r^2 - r)*x + 8*r^3 Compare the x terms we have -6x and 8*( -r^2 - r)*x so -6 = 8(-r^2 - r) solve for r

OpenStudy (whpalmer4):

Note that you'll get two different values for \(r\), and depending on which one you choose, a different value of \(a\). For each value of \(a\), the original equation will have two solutions, and if you do the work correctly, the roots will satisfy the constraint that one is the square of the other. Don't be confused by the fact that the two values of \(r\) you get along the way do not!

OpenStudy (proheiper):

^

zepdrix (zepdrix):

Hmm this was a fun problem :)

OpenStudy (harrycraig):

ITS

zepdrix (zepdrix):

Figure it out Harry Potter? :)

OpenStudy (arnavguddu):

express whole equation in form x^2 + px + q = 0 form (here divide throughout by 8) now if A, B are the roots then A+B = -p AB = q but one root is square of the other so B = A^2.... not solve by this method